A window is 50 cm long. A stone is falling from a height of 40 cm above the window. It crosses the window in

Grade:9

1 Answers

Arun
6 years ago
Dear student
 
Use the kinematic equations as follows:
SUVAT-- s-displacement, u-initial velocity, v-final velocity,a- acceleration, t-timev=u + atv2 = u2 + 2ass=1/2(u+v)ts=ut + 1/2at2

The values for this questions are as follows:
s= 40cm and 50cm
a= 9.8m/s2 (due to gravity)
u=0m/s

The time it takes to reach the top of the window:
s=0.4m
s=ut+1/2at^2
0.4=0xt + 1/2 x 9.8 x t^2
0.4=4.9t^2            t^2= 0.4/4.9
                                 = 0.08163
                                 t= 0.285 s
                                  
The time taken to cross the window:
s=ut+1/2at^2
0.5= (0 x t) + 0.5x9.8xt^2
4.9t^2=0.5 
t^2=0.5/4.9         t^2= 0.1020
                              t=0.1010 seconds

The stone crosses the window in 0.1010 seconds
 
Regards
Arun(askIITians forum expert)